Evaluate (x+y)-1 given x=3 and y=-5
Answer by oscargut(2103)

Evaluate (x+y)-1 given x=3 and y=-5
if x=3 and y=-5 then (x+y)-1=(3+(-5))-1=3-5-1=-3